OpenStudy (anonymous):

Integrate (sqrt x - x)^2 firstly, i expand it then integrate one by one and do the sum together and then add the constant but i am not sure of my answer what do you get?

OpenStudy (accessdenied):

What did you get for your answer?

OpenStudy (anonymous):

x^2/2 - 2/3x^7/2 + x^3/3 +c

OpenStudy (dumbcow):

\[\int\limits_{?}^{?}2x \sqrt{x}dx = 2\int\limits_{?}^{?}x^{3/2}dx = \frac{4}{5}x^{5/2}\]

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

the rest is good

OpenStudy (anonymous):

wut good?

OpenStudy (dumbcow):

the middle term 2xsqrtx

OpenStudy (anonymous):

so my answer is correct or wrong?

OpenStudy (dumbcow):

the middle term is wrong thats what i was showing you the correct integration of the middle term

OpenStudy (anonymous):

okay i get it =)

OpenStudy (anonymous):

so the ansawer is -4/5 x^5/2

OpenStudy (dumbcow):

right

OpenStudy (anonymous):

thanks =d